# Equivalent class of Emergent Single Weyl Fermion in 3d Topological States: gapless superconductors and superfluids Vs chiral fermions

### v1

@@ Abstract
In this article, we put forward a practical but generic approach towards constructing a large family of $(3+1)$ dimension lattice models which can naturally lead to a single Weyl cone in the infrared (IR) limit. Our proposal relies on spontaneous charge $U(1)$ symmetry breaking to evade the usual no-go theorem of a single Weyl cone in a 3d lattice. We have explored three concrete paths in this approach, all involving fermionic topological symmetry protected states (SPTs). Path a) is to push a gapped SPT in a 3d lattice with time-reversal symmetry (or $T$-symmetry) to a gapless topological quantum critical point (tQCP) which involves a minimum change of topologies,i.e. $\delta N_{w}=2$ where $\delta N_{w}$ is the change of winding numbers across the tQCP. Path b) is to peal off excessive degrees of freedom in the gapped SPT via applying $T$-symmetry breaking fields which naturally result in a pair of gapless nodal points of real fermions. Path c) is a hybrid of a) and b) where tQCPs, with $\delta N_{w}\geq 2$, are further subject to time-reversal-symmetry breaking actions. In the infrared limit, all the lattice models with single Weyl fermions studied here are isomorphic to either a tQCP in a DIII class topological superconductor with a protecting $T$-symmetry, or its dual, a $T$-symmetry breaking superconducting nodal point phase, and therefore form an equivalent class. For a generic $T$-symmetric tQCP along Path a), the conserved-charge operators span a six-dimensional linear space while for a $T$-symmetry breaking gapless state along Path b), c), charge operators typically span a two-dimensional linear space instead. Finally, we pinpoint connections between three spatial dimensional lattice chiral fermion models and gapless real fermions that can naturally appear in superfluids or superconductors studied previously.
@@ I Introduction
A very puzzling and distinct feature of tQCPs in SPTs is perhaps the appearance of emergent symmetries, anomalous symmetries that are uniquely related to gapless tQCPs. These anomalous symmetries can not appear in a gapped SPT with the same protecting symmetry and are non-on site ones. It is highly unique, especially in high-dimensions say in 3d lattices.
@@
It is therefore quite natural to discuss potential applications of such unique emergent symmetries in gapless states, either as tQCPs or as stable gapless phases. Our main objective here is to explore in details one such applications, towards an emergent single Weyl fermion in 3d lattices (either left-handed or right-handed) or more precisely lattice chiral fermions via a family of topological gapless states which can be associated to a class of gapped fermonic topological states.
Single Weyl cone dynamics have naturally appeared in quite a few studies of $(3+1)$-dimensional $Z^{T}_{2}$-symmetry protected topological states, either as a representation of infrared dynamics at topological quantum critical points (tQCPs)[33, 41] or as an effective field theory (EFT) of stable nodal point states as a well-known example of gapless topological phases [27, 32, 28]. In Ref.[32], a lattice model was further constructed to verify the paradigm of emergent single Weyl fermion by scanning the whole torus three of the momentum space. As a pleasant surprise, the appearance of such dynamics does not contradict the well-known no-go theorem of single Weyl cones in a three-dimensional (spatial) lattices [42, 43, 44, 45], making it possible to develop a lattice model where the ultraviolet (UV) completion of such an IR Weyl cone can be further explored and understood with fine resolutions.
One of the interesting applications of this observation can be the potential emergence of a single copy of supersymmetric (SUSY) conformal field theory (CFT) in a three-dimensional (or two dimensional) bulk with $N_{f}=\frac{1}{2}$ fermions or one-half of a three dimensional Dirac fermions[33]. Two copies of decoupled SUSY CFTs were originally suggested in a two-dimensional lattice[46]. They were also later suggested to appear in two-dimensional surfaces of three dimensional bulks, where each surface supports one single copy of SUSY CFT[47, 48, 49, 50].
@@
Note that all $3d$ Weyl semi-metals with the charge $U(1)$ symmetry have to have a pair or pairs of Weyl cones[54, 40]. That is fully consistent with the standard fermion doubling theorem[43, 42]. We will be interested in a family of single Weyl fermion lattice models which evade the no-go theorem via breaking the charge $U(1)$ symmetry. As elaborated below, physically they all describe gapless superfluids or superconductors in the infrared limit and therefore natually form an equivalent class.
@@
It is worth mentioning that in $(1+1)D$ a Chiral fermion lattice model can be constructed (via a gapless spinless fermion) with charge $U_{V}(1)$ symmetry unbroken[52]. In addition, the $U_{A}(1)$ is either non-compact or if compact, has an anomaly with $U_{A}(1)$ that is fundamentally non-abelian, i.e. is defined by an Onsager Algebra[52, 58]. Here we will focus exclusively on single Weyl fermion in 3d lattice models.
@@
All the 3d single Weyl fermion lattice models constructed so far that are known to us can be mapped, via $Spin(4)$ unitary transformations, into a $p$-wave superfluid or superconductor. In the infrared limit, they shall be isomorphic to either a gapless superconducting tQCP with the time reversal symmetry or its dual, a superconducting nodal point phase which breaks the time reversal symmetry. Therefore, these models form an equivalent class (see below) and belong to the same family.
3) All the 3d lattice models discussed so far, if they result in single Weyl cone quantum dynamics form an equivalent class of Hamiltonians that are related by Spin(4) unitary transformations. The whole family of lattice Hamiltonians can be further encoded in two dual copies of $3\otimes 3$ dimensional $(1,1)$ representations of a $Spin(4)$ group. And one of the $SU(2)$ subgroups in the $Spin(4)$ group can be identified as a subgroup of emergent Lorentz $SO(3,1)$ group. In addition, the two copies of these theories are further connected by a $\sigma-\tau$ duality transformation previously introduced in Ref.[28].
@@
In this section, we will introduce the two typical paths, Path a) and b), to lattice fermion models both involve gapped gapped fermionic topological symmetry protected states (SPTs) with the time reversal symmetry to start with. In all the discussions below, we only deal with models where the charge $U(1)$ symmetry is broken spontaneously and quantum states or phases in our discussions physically are always gapless superfluids or superconductors. There are either quantum critical points between two topologically distinct SPTs or magnetically strongly polarized SPTs.
@@
Path b): This second path breaks the time reversal symmetry explicitly. We peal off excessive degrees of freedom in a gapped SPT through applying T-symmetry breaking fields to the SPT which naturally result in a pair of gapless real fermion nodal points. A lattice model was previously constructed in ref.[32] to explicitly verify the existence of such a phase where only a pair of real fermion nodal points appear. That study was to verify that only single Weyl cone emerges in the low energy sector near a pair nodal points of real fermion bands[27] in the whole torus-three $\mathcal{T}^{3}$ of the crystal momenta.
In this article, when examining these main paths, we will mainly explore the UV completion of the emergent symmetries associated with the single Weyl cone physics in the infrared limit. We carry out all our discussions via lattice models where the whole momentum space of torus-three $\mathcal{T}^{3}$ can be tracked explicitly.
In Sect.III, we revisit the fermion no-go theorem on band crossings, but carry out our examination on real fermions rather than the usual complex fermions so to apply dircetly to gapless superconductors or superfluids. Instead of using topological homotopy analyses and investigating the topological curvature flux emitted from crossings, here we employ an alternative differential-geometry-based approach.
We introduce a six dimensional manifold that consists of two three dimensionional oriented submanifolds. One of such three-dimensional submanifolds is spanned by the Hamiltonian fibers, $H(\bf p)$ which are momentum $\bf p$} dependent, and the other three dimensional submanifold is $\mathcal{T}^{3}$ for the three dimensional lattice momenta as a base space. Whenever the two oriented submanifolds intersect, we show that two real fermion bands cross each with a specific handedness. We then identify that the numbers of Weyl cones, left or right, can be directly mapped into the numbers of intersects between two oriented three-dimensional manifolds.
We show that for real fermions, the intersections always appear in pairs at $\pm{\bf p}_{i}$,i=1,2,3,Mi=1,2,3...,M. The even-integer numbers of interactions follows a intersect theorem in differential geometry. The momentum space structure of these intersections, i.e. intersections always appear in pairs at $\pm{\bf p}$, is a result of the intrinsic charge conjugation symmetry of real fermions present in any charge $U(1)$ symmetry breaking state.
These intersections therefore indicate $2M$ real fermion band crossings with $M=1,2,3,..$.... The case of $M=1$ has one pair of real fermion band crossings at $\mathbf{p}$ and $\mathbf{p}$ respectively. Because of the charge conjugation symmetry, it can be exactly mapped into an emergent Weyl fermion cone via a straightforward explicit reconstruction. In addition, $M=3,5,..$ lead to odd numbers of Wey fermions.
@@
Model I represents a lattice model of a generic tQCP with protecting time-reversal $T$-symmetry that involved a minimum change of topologies with $\delta N^{f}_{w}=2$. Here $\delta N_{w}$ is the change of topological invariants across a tQCP and the superscript f further refers to the fundamental value. It is a tQCP in the well-known DIII class of topological superconductors with all the UV completed details in torus-three $\mathcal{T}^{3}$.
@@
In Sect.V, we will focus on three lattice models along Path c) which is a hydrid of Path a) and Path b) that involves both tQCPs and time reversal symmetry breaking fields. We start with the model III, a tQCP with a minimum change of topologies $\delta N_{w}=2$ but further subject to a time reversal symmetry breaking magnetic field. The magnetic field lifts the degeneracy of Kramer doublets and results in two real fermion band crossing points at $\pm{\bf p}_{0}$.
@@
In both Sect.IV, V, we also illustrate the UV completion of the emergent symmetries in the infrared limit and construct explicitly the symmetry charge operators in torus-three $\mathcal{T}^{3}$. We further briefly discuss different phases appearing in these lattices models and identify the phase boundaries for the phases with an emergent single Weyl fermion.
In Sect.VI, We then further illustrate how all different models discussed in the previous two sections and known to us can be organized into a distinct linear representation of $Spin(4)$. The infrared (IR) limit of all the lattice models discussed so far, if they lead to single Weyl cone dynamics, shall always be a part of a small family of Hamiltonians that can be related to each other via Spin(4) transformations. Both the lattice Hamiltonians along with their infrared limits form an equivalent class.
@@ II.1 Real fermion representation
One way to avoid the Nielsen-Ninomiya no-go theorem III.1 is to intensionally break the charge U(1) symmetry, the exact conservation of the U(1) charge being the main assumption of the theorem, and the natural first step towards doing so is to consider real fermions.
@@ II.2 Nambu Representation and Real Fermions
whereas in the real fermion basis it comes down to just eq.(Eq.17), as one can verify that $R\tau^{x}R^{T}=iI$.
@@ II.3 Lattice models of real fermions
We will study in details four concrete 3D lattice models in section IV that can be thought of as some particular cases of the general model that we construct below. Starting with a gaped fermionic topological symmetry protected state (SPT) with the T-symmetry, i.e. a DIII class topological superconductor, we can pursue along the two distinct paths a) and b) towards the construction of lattice chiral fermions.
Let us emphasize again here that physically all the 3D lattice model under our considerations lead to gapless superconductors or superfluids where the conventional charge $U(1)$ symmetry is broken spontaneously. They all lead to an emergent single Weyl cone physics in the infrared limit. However, our main interest in this article is about the UV completion of infrared Weyl fermions in the momentum space of $\mathcal{T}^{3}$.
@@
Path a): We push the gaped state to be a quantum critical one while preserving the time reversal symmetry. Especially, we close the gap of our SPT at one of the 8 corners of the Brillouin zone say at at $\mathbf{k}=0$, while maintaining a finite gap everywhere else. The state obtained in this way is a Time reversal invariant three dimensional tQCP.
Path b): we apply a $T$-symmetry breaking field to the fully gaped three dimensional SPT which naturally results in a pair of gapless nodal points. This pair can be further applied to reconstruct a single Weyl cone (Fig2).
@@ II.3.1 Path a)
The coefficient $M({\bf k})$ is a T-symmetry preserving coupling, so the bands are Kramer degenerate everywhere. It can be seen as an effective model of a p-wave superconductor in the limit of strong coupling. As long as this coefficient is non-zero at all of the 8 points $k_{x},k_{y},k_{z}=0,\pi$, the superconducting SPT state remains gaped. We can choose:
@@ II.3.2 Path b)
We start with the previous model in a gaped phase, $\epsilon\neq 0$, and we add a T-breaking field (e.g. a magnetic field along the z-direction). We replaced $3+\epsilon$ by $\mu$ as $\epsilon$ is no longer supposed to go to zero:
@@
Figure 4: (Nodal phase) Spectrum for $\mu=4$ and, left $B=0$, the bands are 2 times degenerate; right $B=6.5$, the magnetic field parameter $B$ lifts the degeneracy and for a range of values gives only two crossings. This allows the pealing off of the excessive degrees of freedom (dotted bands) in this real fermion formalism, a crucial step towards the single Weyl fermion.

### v2

# Equivalence class of Emergent Single Weyl Fermion in 3d Topological States: gapless superconductors and superfluids Vs chiral fermions

@@ Abstract
In this article, we put forward a practical but generic approach towards constructing a large family of $(3+1)$ dimension lattice models which can naturally lead to a single Weyl cone in the infrared (IR) limit. Our proposal relies on spontaneous charge $U(1)$ symmetry breaking via a pair condensate to evade the usual no-go theorem of a single Weyl cone in a 3d lattice. We have explored three concrete paths in this approach, all involving fermionic topological symmetry protected states (SPTs). Path a) is to push a gapped SPT in a 3d lattice with time-reversal symmetry (or $T$-symmetry) to a gapless topological quantum critical point (tQCP) which involves a minimum change of topologies, i.e. $\delta N_{w}=2$ where $\delta N_{w}$ is the change of winding numbers across the tQCP. Path b) is to peel off excessive degrees of freedom in the gapped SPT via applying $T$-symmetry breaking fields which naturally result in a pair of gapless nodal points of real fermions. Path c) is a hybrid of a) and b) where tQCPs, with $\delta N_{w}\geq 2$, are further subject to time-reversal-symmetry breaking actions. We identify detailed phase boundaries of emergent single Weyl fermions in these lattice models. The whole family of lattice models with single Weyl fermions here can be effectively encoded in two dual copies of (1, 1) representations of a Spin(4) group. In the infrared limit, all the lattice models with single Weyl fermions studied here are isomorphic to either a tQCP in a DIII class topological superconductor with a protecting $T$-symmetry, or its dual, a $T$-symmetry breaking superconducting nodal point phase, and therefore form an equivalence class. For a generic $T$-symmetric tQCP along Path a), the conserved-charge operators span a six-dimensional linear space while for a $T$-symmetry breaking gapless state along Path b), c), charge operators typically span a two-dimensional linear space instead. Finally, we pinpoint connections between three spatial dimensional lattice chiral fermion models and gapless real fermions that can naturally appear in superfluids or superconductors studied previously.
@@ I Introduction
A very puzzling and distinct feature of tQCPs in SPTs is perhaps the appearance of emergent symmetries, anomalous symmetries that are uniquely related to gapless tQCPs. These anomalous symmetries can not appear in a gapped SPT with the same protecting symmetry and are non-on site ones. It is highly unique, especially in high-dimensions, e.g. in 3d lattices.
@@
It is therefore quite natural to discuss potential applications of such unique emergent symmetries in gapless states, either as tQCPs or as stable gapless phases. Our main objective here is to explore in details one such application, towards an emergent single Weyl fermion in 3d lattices (either left-handed or right-handed) or more precisely lattice chiral fermions via a family of topological gapless states which can be associated to a class of gapped fermonic topological states.
Single Weyl cone dynamics have naturally appeared in quite a few studies of $(3+1)$-dimensional $Z^{T}_{2}$-symmetry protected topological states, either as a representation of infrared dynamics at topological quantum critical points (tQCPs)[61, 63] or as an effective field theory (EFT) of stable nodal point states as a well-known example of gapless topological phases [58, 28, 62]. In Ref.[28], a lattice model was further constructed to verify the paradigm of emergent single Weyl fermion by scanning the whole 3-torus of the momentum space. As a pleasant surprise, the appearance of such dynamics does not contradict the well-known no-go theorem of single Weyl cones in a three-dimensional (spatial) lattice [38, 39, 19, 29], making it possible to develop a lattice model where the ultraviolet (UV) completion of such an IR Weyl cone can be further explored and understood with fine resolutions.
One of the interesting applications of this observation can be the potential emergence of a single copy of supersymmetric (SUSY) conformal field theory (CFT) in a three-dimensional (or two dimensional) bulk with $N_{f}=\frac{1}{2}$ fermions or one-half of a three dimensional Dirac fermions[61]. The supersymmetri counterpart of two majorana fermion cones here can form a complex fermion under certain conditions, resulting in an emergent $U(1)$ symmetry.
Two copies of decoupled SUSY CFTs were originally suggested in a two-dimensional lattice[33]. They were also later suggested to appear in two-dimensional surfaces of three dimensional bulks, where each surface supports one single copy of SUSY CFT[22, 26, 60, 14].
@@
Note that all $3d$ Weyl semi-metals with the charge $U(1)$ symmetry have to have a pair or pairs of Weyl cones[9, 3]. That is fully consistent with the standard fermion doubling theorem[39, 38]. We will be interested in a family of single Weyl fermion lattice models which evade the no-go theorem via breaking the charge $U(1)$ symmetry. As elaborated below, physically they all describe gapless superfluids or superconductors in the infrared limit and therefore naturally form an equivalence class.
@@
It is worth mentioning that in $(1+1)D$ a Chiral fermion lattice model can be constructed (via a gapless spinless fermion) with charge $U_{V}(1)$ symmetry unbroken[12]. In addition, the $U_{A}(1)$ is either non-compact or if compact, has an anomaly with $U_{A}(1)$ that is fundamentally non-abelian, i.e. is defined by an Onsager Algebra[12, 49]. (See also recent attempts in $(3+1)D$ lattices[10].) Here we will focus exclusively on single Weyl fermion in 3d lattice models.
@@
All the 3d single Weyl fermion lattice models constructed so far that are known to us can be mapped, via $Spin(4)$ unitary transformations, into a $p$-wave superfluid or superconductor. In the infrared limit, they shall be isomorphic to either a gapless superconducting tQCP with the time reversal symmetry or its dual, a superconducting nodal point phase which breaks the time reversal symmetry. Therefore, these models form an equivalence class (see below) and belong to the same family.
3) All the 3d lattice models discussed so far, if they result in single Weyl cone quantum dynamics form an equivalence class of Hamiltonians that are related by Spin(4) unitary transformations. The whole family of lattice Hamiltonians can be further encoded in two dual copies of $3\otimes 3$ dimensional $(1,1)$ representations of a $Spin(4)$ group. And one of the $SU(2)$ subgroups in the $Spin(4)$ group can be identified as a subgroup of emergent Lorentz $SO(3,1)$ group. In addition, the two copies of these theories are further connected by a $\sigma-\tau$ duality transformation previously introduced in Ref.[62].
@@
In this section, we will introduce the two typical paths, Path a) and b), to lattice fermion models both involve gapped fermionic topological symmetry protected states (SPTs) with the time reversal symmetry to start with. In all the discussions below, we only deal with models where the charge $U(1)$ symmetry is broken spontaneously and quantum states or phases in our discussions physically are always gapless superfluids or superconductors. There are either quantum critical points between two topologically distinct SPTs or magnetically strongly polarized SPTs.
@@
Path b): This second path breaks the time reversal symmetry explicitly. We peel off excessive degrees of freedom in a gapped SPT through applying T-symmetry breaking fields to the SPT which naturally result in a pair of gapless real fermion nodal points. A lattice model was previously constructed in ref.[28] to explicitly verify the existence of such a phase where only a pair of real fermion nodal points appear. That study was to verify that only single Weyl cone emerges in the low energy sector near a pair nodal points of real fermion bands[58] in the whole 3-torus $\mathbb{T}^{3}$ of the crystal momenta.
In this article, when examining these main paths, we will mainly explore the UV completion of the emergent symmetries associated with the single Weyl cone physics in the infrared limit. We carry out all our discussions via lattice models where the whole momentum space of 3-torus $\mathbb{T}^{3}$ can be tracked explicitly.
In Sect.III, we revisit the fermion no-go theorem on band crossings, but carry out our examination on real fermions rather than the usual complex fermions so to apply directly to gapless superconductors or superfluids. Instead of using topological homotopy analyses and investigating the topological curvature flux emitted from crossings, here we employ an alternative differential-geometry-based approach.
We introduce a six dimensional manifold that consists of two three-dimensional oriented submanifolds. One of such three-dimensional submanifolds is spanned by the Hamiltonian fibers, $H(\bf p)$ which are momentum $\bf p$} dependent, and the other three dimensional submanifold is $\mathbb{T}^{3}$ for the three dimensional lattice momenta as a base space. Whenever the two oriented submanifolds intersect, we show that two real fermion bands cross each with a specific handedness. We then identify that the numbers of Weyl cones, left or right, can be directly mapped into the numbers of intersections between two oriented three-dimensional manifolds.
We show that for real fermions, the intersections always appear in pairs at $\pm{\bf p}_{i}$,i=1,2,3,Mi=1,2,3...,M. The even-integer numbers of interactions follows an intersection theorem in differential geometry. The momentum space structure of these intersections, i.e. intersections always appear in pairs at $\pm{\bf p}$, is a result of the intrinsic charge conjugation symmetry of real fermions present in any charge $U(1)$ symmetry breaking state.
These intersections therefore indicate $2M$ real fermion band crossings with $M=1,2,3,..$.... The case of $M=1$ has one pair of real fermion band crossings at $\mathbf{p}$ and $\mathbf{p}$ respectively. Because of the charge conjugation symmetry, it can be exactly mapped into an emergent Weyl fermion cone via a straightforward explicit reconstruction. In addition, $M=3,5,..$ lead to odd numbers of Weyl fermions.
@@
Model I represents a lattice model of a generic tQCP with protecting time-reversal $T$-symmetry that involved a minimum change of topologies with $\delta N^{f}_{w}=2$. Here $\delta N_{w}$ is the change of topological invariants across a tQCP and the superscript f further refers to the fundamental value. It is a tQCP in the well-known DIII class of topological superconductors with all the UV completed details in 3-torus $\mathbb{T}^{3}$.
@@
In Sect.V, we will focus on three lattice models along Path c) which is a hybrid of Path a) and Path b) that involves both tQCPs and time reversal symmetry breaking fields. We start with the model III, a tQCP with a minimum change of topologies $\delta N_{w}=2$ but further subject to a time reversal symmetry breaking magnetic field. The magnetic field lifts the degeneracy of Kramer doublets and results in two real fermion band crossing points at $\pm{\bf p}_{0}$.
@@
In both Sect.IV, V, we also illustrate the UV completion of the emergent symmetries in the infrared limit and construct explicitly the symmetry charge operators in torus-three $\mathbb{T}^{3}$. We further briefly discuss different phases appearing in these lattices models and identify the phase boundaries for the phases with an emergent single Weyl fermion.
In Sect.VI, We then further illustrate how all different models discussed in the previous two sections and known to us can be organized into a distinct linear representation of $Spin(4)$. The infrared (IR) limit of all the lattice models discussed so far, if they lead to single Weyl cone dynamics, shall always be a part of a small family of Hamiltonians that can be related to each other via Spin(4) transformations. Both the lattice Hamiltonians along with their infrared limits form an equivalence class.
@@ II.1 Real fermion representation
One way to avoid the Nielsen-Ninomiya no-go theorem III.1 is to break the charge U(1) symmetry, the exact conservation of the U(1) charge being the main assumption of the theorem, and the natural first step towards doing so is to consider real fermions.
@@ II.2 Nambu Representation and Real Fermions
whereas in the real fermion basis it comes down to just (Eq.17), as one can verify that $R\tau^{x}R^{T}=iI$.
@@ II.3 Lattice models of real fermions
We will study in details four concrete 3D lattice models in section IV that can be thought of as some particular cases of the general model that we construct below. Starting with a gapped fermionic topological symmetry protected state (SPT) with the T-symmetry, i.e. a DIII class topological superconductor, we can pursue along the two distinct paths a) and b) towards the construction of lattice chiral fermions.
Let us emphasize again here that physically all the 3D lattice model under our considerations lead to gapless superconductors or superfluids where the conventional charge $U(1)$ symmetry is broken spontaneously. They all lead to an emergent single Weyl cone physics in the infrared limit. However, our main interest in this article is about the UV completion of infrared Weyl fermions in the momentum space of $\mathbb{T}^{3}$.
@@
Path a): We push the gapped state to be a quantum critical one while preserving the time reversal symmetry. Especially, we close the gap of our SPT at one of the 8 corners of the Brillouin zone say at at $\mathbf{k}=0$, while maintaining a finite gap everywhere else. The state obtained in this way is a Time reversal invariant three dimensional tQCP.
Path b): we apply a $T$-symmetry breaking field to the fully gapped three dimensional SPT which naturally results in a pair of gapless nodal points. This pair can be further applied to reconstruct a single Weyl cone (Fig2).
@@ II.3.1 Path a)
The coefficient $M({\bf k})$ is a T-symmetry preserving coupling, so the bands are Kramer degenerate everywhere. It can be seen as an effective model of a p-wave superconductor in the limit of strong coupling. As long as this coefficient is non-zero at all of the 8 points $k_{x},k_{y},k_{z}=0,\pi$, the superconducting SPT state remains gapped. We can choose:
@@ II.3.2 Path b)
We start with the previous model in a gapped phase, $\epsilon\neq 0$, and we add a T-breaking field (e.g. a magnetic field along the z-direction). We replaced $3+\epsilon$ by $\mu$ as $\epsilon$ is no longer supposed to go to zero. For the purpose of the current study, we consider the magnetic field $B$ to be k-independent:
@@
Figure 4: (Nodal phase) Spectrum for $\mu=2.5$ and, left $B=0$, the bands are 2 times degenerate; right $B=4$, the magnetic field parameter $B$ lifts the degeneracy and for a range of values gives only two crossings. This allows the peeling off of the excessive degrees of freedom (dotted bands) in this real fermion formalism, a crucial step towards the single Weyl fermion.